OpenStudy (anonymous):

Simplify the expression. Write the answer without using parentheses or negative exponents (ab^3)^-4

OpenStudy (anonymous):

\[\left( ab^3 \right)^{-4}=\frac{ 1 }{ \left( ab^3 \right)^4 }=?\]

OpenStudy (anonymous):

So, what do I need to dothere.

OpenStudy (anonymous):

\[\left( x^2y^3 \right)^5=x ^{2*5}y ^{3*5}=x ^{10}y ^{15}\] a has exponent 1,b has exponent 3, can you solve? i have given you an example.

OpenStudy (anonymous):

you got it or should i proceed.

OpenStudy (anonymous):

\[=\frac{ 1 }{ a^4b ^{3*4} }=\frac{ 1 }{ a^4b ^{12} }\]

OpenStudy (anonymous):

Yeah is that the answer than ?

OpenStudy (anonymous):

@surjithayer